byeeee:

What is the equation for an arithmetic sequence with a first term of 6 and a second term of 2? an = 6 − 2(n − 1) an = 6 + 2(n − 1) an = 6 + 4(n − 1) an = 6 − 4(n − 1)

byeeee:

@dude

dude:

Haven't done much arithmetic lately but do you have an idea on this?

byeeee:

well i have the formula and i know the 1st term is 6 but i don't know about the second term

dude:

So for 6 to go to 2, how much would you need to increase/decrease by?

byeeee:

4

dude:

 This is the general equation

dude:

In this case it is re-written as \(a_n=a_1+d(n-1)\) Which is the same thing, we know that \(a_1=6\), so \(a_n=6+d(n-1)\) If you have to go down 4, then the rate is simply -4 \(a_n=6-4(n-1)\)

byeeee:

ohhh

dude:

If you want to check \(a_n=6-4(n-1)\) Substitute: (n=2 To find the second number) \(a_2=6-4(2-1)\) \(a_2=6-4(1)\\ a_2=6-4\\ a_2=2\)

byeeee:

so you have to actually put the -4 into the equation

dude:

Yes

byeeee:

thank you!